# Stabilizability and lower spectral radius for linear switched systems with singular matrices

### v2

@@ Abstract
We investigate the stabilizability of linear discrete-time switched systems with singular matrices, focusing on the spectral radius in this context. A new lower bound of the stabilizability radius is proposed, which is applicable to any matrix set. Based on this lower bound, more relationships between the stabilizability radius and joint spectral subradius are established. Detailed analysis of the stabilizability radius of a special kind of two-dimensional switched system, consisting of a singular matrix and a matrix with complex eigenvalues, is presented. The Hausdorff dimensions of the parameter sets such that the stabilizability radius of these systems equals a constant are also presented. Other properties of switched systems with singular matrices are also discussed along with examples.
@@ 1 Introduction
The switched system is widely studied in theory and applications. Let us mention [19] and [20] for comprehensive descriptions of formation and stability of such systems. Ref. [3] contains more recent works on continuous time switched linear systems. There are many aspects of research on switched systems. For example, Ref. [16] consider the control of switched linear systems, the Lyapunov analysis of switched systems considered in [4] and [5] and Ref. [2] is about applications of switched systems to epidemiological spread.
Despite the simple form of the system itself, the stability and stabilizability of such systems are hard to determine, even for linear ones. The discrete-time linear switched system of the given matrix set $\mathcal{M}\subseteq\mathbb{R}^{n\times n},\;n\geq 2$ is given by
@@
where $M_{\sigma(t)}\in\mathcal{M}$, $\sigma(t)\in\{1,\dots,m\}$ and $m$ is the cardinality of $\mathcal{M}$. There are many general results on the stability and stabilizability of such system (for example, see [20]). However, we are not aware of any study focusing on the case where the matrix set $\mathcal{M}$ contains singular matrices. In the literature, the term “singular switched system” refers to a special kind of switched system with many known results (see [21] and [13]), which is different from the one we consider in this paper.
To describe the stability and stabilizability of system (1.1), the joint spectral properties are widely used. Let $\mathcal{M}^{T}$ denote the set that contains all $M_{\sigma(t)}\cdots M_{\sigma(1)}$, the joint spectral radius of system (1.1) is defined as
@@
for any norm. The joint spectral subradius is defined as
@@
$\tilde{\rho}=\sup_{x_{0}\in\mathbb{R}^{n}}\inf\{\lambda\geq 0\,|\,\exists\sigma(\cdot),c>0\;s.t.\;\|x_{\sigma}(t)\|\leq c\lambda^{t}\|x_{0}\|\,\forall t\geq 0\},$ (1.4)
@@
  • A switched system is called stable if $\hat{\rho}(\mathcal{M})<1$ (see, for example, [9]).
@@
The stabilizability radius $\tilde{\rho}$, first introduced in [11], had many implicit studies beforehand, for example, Ref. [12] considered this problem in measurement scheduling and other cases, and Ref. [7] considered this problem with applications in HIV drug control. For non-singular matrix sets, $\tilde{\rho}$ has been proved to have several lower bounds in [6].
@@
In applications mentioned above, the matrix set could contain singular matrices. In this paper, we have a new lower bound of the stabilizability radius for any finite matrix set. In the meantime, we do a detailed analysis of $\tilde{\rho}$ of a two dimensional switched system with singular matrices, and show how to compute the stability radius $\tilde{\rho}$ using continued fractions and Diophantine approximation. We also show that for the case $\tilde{\rho}=0$, the set of such $\mathcal{M}$ has zero Hausdorff dimension. Other properties of systems are mentioned as they appear.
@@
Some notations are used in this paper with their well-known meanings. The area denotes the n1n-1 dimensional Lebesgue measure of a set in $\mathbb{R}^{n}$. For two functions $f$ and $g$, we write $f(x)\sim g(x)$ as $x\rightarrow 0$ if $\lim_{x\rightarrow 0}f(x)/g(x)=1$. The notation $\|\cdot\|$ will now specifically denote the Euclidean norm in $\mathbb{R}^{n}$.
@@
In section $2$, we show the $\tilde{\rho}$ is lower bounded by $\check{\rho}$ divided by the cardinality of $\mathcal{M}$ and give one more situation when $\tilde{\rho}=\check{\rho}$. Section $3$ is dedicated to the stabilizability radius of the two dimensional switched system with a singular matrix and a matrix with complex eigenvalues. We also show the Hausdorff dimension of the parameter set such that $\tilde{\rho}=c,0\leq c<1$ is zero. Section $4$ contains several examples of switched systems with singular matrices about relationship between smallest singular value and $\tilde{\rho}$. One strategy of finding the optimal switching law is also mentioned.
2 Lower Bound
2.1 Lower bound formation
@@
Let $A_{t}$ denote the product $M_{\sigma(t)}\cdots M_{\sigma(1)}$. Suppose $s^{2}_{1}(A_{t}),\dots,s^{2}_{n}(A_{t})$ are ordered eigenvalues of $A^{T}_{t}A_{t}$ and $v_{1},\dots,v_{n}$ are corresponding orthonormal eigenvectors, where $s_{1}$ is smallest singular value of $A_{t}$. We have two nice properties of $v_{i}$.
@@
    $\|A_{t}v_{i}\|=s_{i}(A_{t})$.
Suppose $s_{1},\dots,s_{\kappa}$ are zero singular values, then $v_{1},\dots,v_{\kappa}$ form an orthonormal basis for $\mathbf{Ker}(A_{t})$. Let $(x_{1},\dots,x_{n})$ denote the coordinates of $x\in\mathbb{R}^{n}$ in terms of $v_{1},\dots,v_{n}$.
Start from the unit sphere. If the $t$ iteration of the unit sphere under any $A_{t}$ is a subset of a ball centered at the origin, we know that $\|x(t)\|\leq\lambda\|x_{0}\|$, where $\lambda$ is the radius of the ball. Let $B_{r}$ denote the ball with radius $r$ centered at the origin and $S^{n-1}$ denote the $n$ dimensional unit sphere. We consider the intersection of the pre-image of $A_{t}$ and $S^{n-1}$, i.e. $S_{r,A_{t}}=\{x\in S^{n-1}|\|A_{t}x\|\leq r\}$. As $A_{t}v_{j}=0$ for $j=1,\dots,\kappa$, then for any $x\in S_{r,A_{t}}$, we have
$\|A_{t}x\|^{2}=\|\sum_{i=1}^{n}x_{i}A_{t}v_{j}\|^{2}=\sum_{i=\kappa+1}^{n}s^{2}_{i}(A_{t})x_{i}^{2}\leq r^{2}.$ (2.1)
For any nonzero $A_{t}$, consider the set $S^{\prime}=\{x\in S^{n-1}||x_{n}|\leq r/s_{n}(A_{t})\}$. Since
$\sum_{i=\kappa+1}^{n}s^{2}_{i}(A_{t})x_{i}^{2}\geq s^{2}_{n}(A_{t})x_{n}^{2},$ (2.2)
we have $S_{r,A_{t}}\subseteq S^{\prime}$, and $S^{\prime}$ is a spherical segment. The area of $S^{\prime}$ can be obtained using the area of spherical caps. A formula for area of caps is given by Lemma 2.1.
@@
((1) in [14])The surface area of the cap of the unit sphere $S^{n-1}$, i.e. $T_{n}(h)=\{x\in S^{n-1}|x_{1}\geq h\}$ with $x_{1}$ be one of coordinates of $x$, is
@@
One can prove this by changing variables in the integral definition of $I(h;\cdot,\cdot)$.
@@
The area of $S_{r,A_{t}}$ is upper bounded by
$|S_{r,A_{t}}|\leq|S^{n-1}|I\left(\frac{r^{2}}{s^{2}_{n}\left(A_{t}\right)};\frac{1}{2},\frac{n-1}{2}\right).$ (2.6)
@@
$|S^{\prime}|=|S^{n-1}|-2\left|T_{n}\left(\frac{r}{s_{n}\left(A_{t}\right)}\right)\right|.$ (2.7)
Hence
$|S_{r,A_{t}}|\leq|S^{n-1}|\left(1-I\left(1-\frac{r^{2}}{s^{2}_{n}\left(A_{t}\right)};\frac{n-1}{2},\frac{1}{2}\right)\right)=|S^{n-1}|I\left(\frac{r^{2}}{s^{2}_{n}\left(A_{t}\right)};\frac{1}{2},\frac{n-1}{2}\right).$ (2.8)
Thus we have a bound of $|S_{r,A_{t}}|$. ∎
Theorem 1.
The stabilizability radius satisfies
@@
The $\tilde{\rho}\leq\check{\rho}$ is from definition of $\tilde{\rho}$. For a given $T$, the area of the union of $S_{r,A_{T}}$ under all $A_{T}$ have an upper bound given by
$\left|\cup_{\sigma}S_{r,A_{T}}\right|\leq\sum_{\sigma}|S_{r,A_{T}}|\leq\sum_{\sigma}|S^{n-1}|I\left(\frac{r^{2}}{s^{2}_{n}(A_{T})};\frac{1}{2},\frac{n-1}{2}\right)$ (2.10)
The sum above is over all possible value of $\sigma(t)$ for $t=1,\dots,T$. Since I(h,1/2,(n1)/2)I(h,1/2,(n-1)/2) is the cumulative distribution function of the beta distribution, I(h,1/2,(n1)/2)I(h,1/2,(n-1)/2) monotonically increases for $0\leq h\leq 1$. Let $S(T)$ denotes the smallest $s_{n}(A_{T})$ for all $\sigma$, then we have
$I\left(\frac{r^{2}}{s^{2}_{n}(A_{T})};\frac{1}{2},\frac{n-1}{2}\right)\leq I\left(\frac{r^{2}}{S(T)^{2}};\frac{1}{2},\frac{n-1}{2}\right).$ (2.11)
Then
$\left|\bigcup_{\sigma}S_{r,A_{T}}\right|\leq m^{T}|S^{n-1}|I\left(\frac{r^{2}}{S(T)^{2}};\frac{1}{2},\frac{n-1}{2}\right).$ (2.12)
From the definition of $\tilde{\rho}$, for any $\lambda>\tilde{\rho}$, there exists a positive constant $c$ such that the union of $S_{c\lambda^{T},A_{T}}$ covers the unit sphere. Hence
$|S^{n-1}|=\left|\bigcup_{\sigma}S_{c\lambda^{T},A_{T}}\right|\leq m^{T}|S^{n-1}|I\left(\frac{c^{2}\lambda^{2T}}{S(T)^{2}};\frac{1}{2},\frac{n-1}{2}\right).$ (2.13)
@@
Consider the limit
$\lim_{T\rightarrow\infty}S(T)^{1/T}.$ (2.15)
An alternative way to write this limit is
$\liminf_{T\rightarrow\infty}\{s_{n}(A_{T})^{1/T}\;|\;\sigma(t)\in\{1,\dots,m\},t\in\{1,\dots,T\}\}.$ (2.16)
Then this is the $\check{\rho}$ of the matrix set $\mathcal{M}$ (see for example Theorem 1.1 in [9]).
Now, we know the asymptotic behavior of I(h;1/2,(n1)/2)I(h;1/2,(n-1)/2) when $h$ tends to 0, allowing us to obtain a lower bound of $\tilde{\rho}$ in terms of $\check{\rho}$. Notice that
@@
for a constant $C$. Let $g(x)$ denote the integrand, then for any $n\geq 2$, we have as $x\rightarrow 0$,
$g(x)\sim x^{-\frac{1}{2}}.$ (2.18)
Namely,
@@
$2C\frac{c\lambda^{T}}{S(T)}\geq\frac{1}{m^{T}},$ (2.20)
which is
$(2Cc)^{1/T}\lambda\geq\frac{S(T)^{1/T}}{m}.$ (2.21)
Let $T$ tend to infinity, then we have for any $\lambda>\tilde{\rho}$,
$\lambda\geq\frac{\check{\rho}}{m}.$ (2.22)
Let $\lambda$ tend to $\tilde{\rho}$, then
$\tilde{\rho}\geq\frac{\check{\rho}}{m}.$ (2.23)
In general this is not an accurate bound, especially for large $m$.
Example 2.3.
For Example 2.3, we have $\check{\rho}=1$. From Theorem 1, we know that $\tilde{\rho}\geq 1/2$. However, the lower bound given by Theorem $2$ in [6] already gives us $\tilde{\rho}\geq\sqrt{2}/2$. The advantage of the bound in Theorem 1 is that it works for matrix sets with singular matrices, especially for those system with known $\check{\rho}$, while lower bounds in [6] doesn’t work for such matrix sets.
Example 2.4.
Switched system with $\mathcal{M}=\{M_{1},M_{2}\}$ such that
In Example 2.4, $M_{1}$ projects any three dimensional vector onto the $(x_{1},x_{2})$ plane. The $\check{\rho}(\mathcal{M})$ of this system equals the $\check{\rho}$ of Example 2.3. As $M_{1}$ is singular, any previous lower bound in the literature doesn’t work for this system. From Theorem 1, we obtain the nontrivial bound $\tilde{\rho}\geq 1/2$.
2.2 Joint spectral subradius and stabilizability radius
From Theorem 1, we know that the stabilizability radius $\tilde{\rho}$ is equivalent to the joint spectral subradius $\check{\rho}$. An immediate corollary of Theorem 1 is that when $\tilde{\rho}=0$, we have $\check{\rho}=0$. This describes the case when a matrix set pointwise converges to zero quickly enough, we know it converges to zero uniformly.
Notice that for a given $\mathcal{M}$, the optimal switching law of $\check{\rho}$ and $\tilde{\rho}$ not necessarily contain every matrix in $\mathcal{M}$. To avoid such phenomena, we have following definition.

### v3

@@ Abstract
We investigate the stabilizability of linear discrete-time switched systems with singular matrices, focusing on the spectral radius in this context. A new lower bound of the stabilizability radius is proposed, which is applicable to any matrix set. Switched systems with rank one singular matrices are discussed: The stabilizability radius and the joint spectral subradius are equal for such systems. Detailed analysis of the stabilizability radius of two-dimensional switched systems, consisting of a singular matrix and a matrix with complex eigenvalues or real eigenvalues, are presented. The condition when an infinitely long aperiodic optimal sequence appears of such system is also discussed. Other properties of switched systems with singular matrices are also discussed along with examples.
@@ 1 Introduction
Switched systems have been widely studied in both theoretical and applied contexts. Comprehensive descriptions of the formation and stability of such systems can be found in [22] and [23]. Ref. [3] contains more recent works on continuous time switched linear systems. From a control and analytical perspective, Ref. [17] consider the control of switched linear systems, the Lyapunov analysis of switched systems considered in [4] and [5]. Beyond purely theoretical inquiries, switched systems have also proven effective in modeling real-world phenomena, such as the epidemiological spread framework proposed by [2] and designing of multiphase trajectories for aerospace vehicles in [19].
Despite their structurally simple form, determining the stability and stabilizability of these systems remains highly challenging, even in the linear case. The discrete-time linear switched system of the given matrix set $\mathcal{M}\subseteq\mathbb{R}^{n\times n},\;n\geq 2$ is given by
@@
where $M_{\sigma(t)}\in\mathcal{M}$, $\sigma(t)\in\{1,\dots,m\}$ and $m$ is the cardinality of $\mathcal{M}$. Notice that $m$ may be infinite. However, by Corollary $3.4$ and Theorem $3.5$ in [20], it is sufficient to only consider finite $m$ in this paper. There are many general results on the stability and stabilizability of such system (for example, see [23]). However, we are not aware of any study focusing on the case where the matrix set $\mathcal{M}$ contains singular matrices. In the literature, the term “singular switched system” refers to a special kind of switched system with many known results (see [24] and [14]), which is different from the one we consider in this paper.
To describe the stability and stabilizability of system (1.1), the joint spectral properties are widely used. Let $\mathcal{M}^{T}$ denote the set that contains all $M_{\sigma(t)}\cdots M_{\sigma(1)}$. Then the joint spectral radius of system (1.1) is defined as
@@
for any norm. Then the joint spectral subradius is defined as
@@
$\tilde{\rho}=\sup_{x_{0}\in\mathbb{R}^{n}}\inf\{\lambda\geq 0\,|\,\exists\sigma(\cdot),c>0\;s.t.\;\|x(t)\|\leq c\lambda^{t}\|x_{0}\|\,\forall t\geq 0\},$ (1.4)
@@
  • A switched system is called asymptotically stable if and only if $\hat{\rho}(\mathcal{M})<1$ (see, for example, [7]).
@@
The stabilizability radius $\tilde{\rho}$, first introduced in [11], had many implicit studies beforehand, for example, Ref. [13] considered this problem in measurement scheduling and other cases, and Ref. [8] considered this problem with applications in HIV drug control. For non-singular matrix sets, $\tilde{\rho}$ has been proved to have several lower bounds in [6], but it would be good to know a lower bound of $\tilde{\rho}$ for an arbitrary matrix set. Let us mention two simple examples
Example 1.1.
Example 1.2.
Switched system with $\mathcal{M}=\{M_{1},M_{2}\}$ such that
This Example 1.1 has been well studied in [6]; the lower bound given by Theorem $2$ in [6] gives us $\tilde{\rho}\geq\sqrt{2}/2$ for Example 1.1. In Example 1.2, $M_{1}$ projects any three dimensional vector onto the $(x_{1},x_{2})$ plane, thus Example 1.1 and 1.2 should have the same dynamics. However, as far as we know, none of the existing lower bound works for Example 1.2, because $M_{1}$ is a singular matrix. Also in applications mentioned above, the matrix set could contain singular matrices.
This lead us to two natural questions:
  • Is there a more general lower bound that remains valid in the presence of singular matrices?
  • What is the structure of linear switched systems with singular matrices, and in particular, how does singularity affect their stabilizability?
By addressing these two questions, we aim to gain a deeper understanding of linear switched systems with singular matrices.
@@
In this paper, we establish a new lower bound of the stabilizability radius for any finite matrix set. In the meantime, we do a detailed analysis of $\tilde{\rho}$ of a two dimensional switched system with singular matrices, and show how to compute the stability radius $\tilde{\rho}$ using continued fractions and Diophantine approximation. We also investigate the finiteness property (formally defined in Definition 4.8), i.e. having finite or periodic optimal sequence, of these two dimensional cases. Other properties of systems are mentioned as they appear.
@@
Some notations are used in this paper with their well-known meanings. The area denotes the n1n-1 dimensional Lebesgue measure of a set in $\mathbb{R}^{n}$. For two functions $f$ and $g$, we write $f(x)\sim g(x)$ as $x\rightarrow 0$ if $\lim_{x\rightarrow 0}f(x)/g(x)=1$. The notation $\|\cdot\|$ will now specifically denote the Euclidean norm in $\mathbb{R}^{n}$. The notation $\|\cdot\|_{N}$ from number theory denotes the distance to the closed integer. $\mathbb{N}$ denotes natural numbers including zero.
@@
In section $2$, we show the $\tilde{\rho}$ is lower bounded by $\check{\rho}$ divided by the cardinality of $\mathcal{M}$. Section $3$ is dedicated to the switched systems with rank one singular matrices, The lower spectral radius equals stabilizability radius for such systems. Other properties of such systems are discussed as well. In section $4$ and $5$, we discuss the stabilizability radius of the two dimensional switched system with two matrices and one of them is a singular matrix. Section $4$ is about switched system with a singular matrix and a rotation matrix, i.e. with complex eigenvalues. We also show the parameter set, such that $\tilde{\rho}=c,0\leq c<1$ and the optimal sequence is infinitely long and aperiodic, has zero Hausdorff dimension in this section. Section $5$ is about switched system with a singular matrix and a hyperbolic matrix, i.e. with real eigenvalues. The condition when optimal sequence is infinitely long and aperiodic of these systems is also discussed. Section $6$ contains several examples of switched systems with singular matrices about the relationship between the smallest singular value and $\tilde{\rho}$. One strategy of finding the optimal switching law is also mentioned.
2 Lower Bound for the stabilizability radius
@@
This section pertains to arbitrary sets of matrices, singular or non-singular. Let $A_{\sigma,t}$ denote the product $M_{\sigma(t)}\cdots M_{\sigma(1)}$. Suppose $s^{2}_{1}(A_{\sigma,t}),\dots,s^{2}_{n}(A_{\sigma,t})$ are ordered eigenvalues of $A^{T}_{\sigma,t}A_{\sigma,t}$ and $v_{1},\dots,v_{n}$ are corresponding orthonormal eigenvectors, where $s_{1}$ is smallest singular value of $A_{\sigma,t}$. We have two properties of $v_{i}$.
Lemma 2.1.
The $v_{i}$ for $i=1,\dots,n$, called singular vectors, satisfies
@@
    $\|A_{\sigma,t}v_{i}\|=s_{i}(A_{\sigma,t})$.
for $i=1,\dots,n$.
Proof.
For the first property, we have
$\left<A_{\sigma,t}v_{i},A_{\sigma,t}v_{j}\right>=\left<v_{i},A^{T}_{\sigma,t}A_{\sigma,t}v_{j}\right>=s^{2}_{j}\left<v_{i},v_{j}\right$>. (2.1)
Then $A_{\sigma,t}v_{i}$ are orthogonal because $v_{i}$ are orthogonal, for $i=1,\dots,n$. For the second property, we have
$\|A_{\sigma,t}v_{i}\|^{2}=\left<A_{\sigma,t}v_{i},A_{\sigma,t},v_{i}\right>=\left<v_{i},A^{T}_{\sigma,t}A_{\sigma,t}v_{i}\right>=s^{2}_{i}.$ (2.2)
The second property follows from taking square root of both sides of (2.2). ∎
Let $B_{r}$ denote the ball with radius $r$ centered at the origin and $S^{n-1}$ denote the $n$ dimensional unit sphere. For $x_{0}\in S^{n-1}$, we have $x(t)\in B_{r}$ for some $r$. Now we want a cover of the intersection of the pre-image of $A_{\sigma,t}$ and $S^{n-1}$, i.e. $S_{r,A_{\sigma,t}}=\{x\in S^{n-1}|\|A_{\sigma,t}x\|\leq r\}$. Suppose $s_{1},\dots,s_{\kappa}$ are zero singular values, then $v_{1},\dots,v_{\kappa}$ form an orthonormal basis of $\mathbf{Ker}(A_{\sigma,t})$. Let $(x_{1},\dots,x_{n})$ denote the coordinate of $x\in\mathbb{R}^{n}$ in terms of $v_{1},\dots,v_{n}$. We can characterize $S_{r,A_{\sigma}}$ as follows:
@@
For any nonzero $A_{\sigma,t}$, let $S^{\prime}=\{x\in S^{n-1}||x_{n}|\leq r/s_{n}(A_{\sigma,t})\}$, then $S_{r,A_{\sigma,t}}\subseteq S^{\prime}$
Proof.
Since $A_{\sigma,t}v_{j}=0$ for $j=1,\dots,\kappa$, for any $x\in S_{r,A_{\sigma,t}}$, we have
$\|A_{\sigma,t}x\|^{2}=\|\sum_{i=1}^{n}x_{i}A_{\sigma,t}v_{j}\|^{2}=\sum_{i=\kappa+1}^{n}s^{2}_{i}(A_{\sigma,t})x_{i}^{2}\leq r^{2}.$ (2.3)
Notice that
$\sum_{i=\kappa+1}^{n}s^{2}_{i}(A_{\sigma,t})x_{i}^{2}\geq s^{2}_{n}(A_{\sigma,t})x_{n}^{2}.$ (2.4)
We have $S_{A_{\sigma,t},r}\subseteq S^{\prime}$. ∎
This $S^{\prime}$ in Lemma 2.2 is a spherical segment. The area of $S^{\prime}$ can be obtained using the area of spherical caps. A formula for area of caps is given by Lemma 2.3.
Lemma 2.3.
((1) in [15])The surface area of the cap of the unit sphere $S^{n-1}$, i.e. $T_{n}(h)=\{x\in S^{n-1}|x_{1}\geq h\}$ with $x_{1}$ be one of coordinates of $x$ and $h>0$, is
@@
One can prove this by changing variables in the integral definition of $I(h;\cdot,\cdot)$. Now by Lemmas 2.2 and 2.3, we can generate a upper bound of $|S_{A_{\sigma,t},r}|$.
@@
The area of $S_{r,A_{\sigma,t}}$ is upper bounded by
$|S_{r,A_{\sigma,t}}|\leq|S^{n-1}|I\left(\frac{r^{2}}{s^{2}_{n}\left(A_{\sigma,t}\right)};\frac{1}{2},\frac{n-1}{2}\right).$ (2.8)
@@
$|S^{\prime}|=|S^{n-1}|-2\left|T_{n}\left(\frac{r}{s_{n}\left(A_{\sigma,t}\right)}\right)\right|.$ (2.9)
By Lemma 2.2, we have
$|S_{r,A_{\sigma,t}}|\leq|S^{n-1}|\left(1-I\left(1-\frac{r^{2}}{s^{2}_{n}\left(A_{\sigma,t}\right)};\frac{n-1}{2},\frac{1}{2}\right)\right)=|S^{n-1}|I\left(\frac{r^{2}}{s^{2}_{n}\left(A_{\sigma,t}\right)};\frac{1}{2},\frac{n-1}{2}\right).$ (2.10)
This upper bound of the area of $S_{r,A_{\sigma,t}}$ leads to a lower bound of stabilizability radius, but we need one more lemma to prove it.
Lemma 2.5.
For an integer $T\geq 1$, let $S(T)$ denotes the smallest $s_{n}(A_{\sigma,T})$ for all $\sigma(\cdot)$. We have
$\liminf_{T\rightarrow\infty}S(T)^{1/T}=\liminf_{T\rightarrow\infty}\{s_{n}(A_{\sigma,t})^{1/t}\;|\;\sigma(t)\in\{1,\dots,m\},t\in\{1,\dots,T\}\},$ (2.11)
i.e.
$\liminf_{T\rightarrow\infty}S(T)^{1/T}=\check{\rho}.$ (2.12)
Proof.
Clearly, $S(T)\in\{s_{n}(A_{\sigma,t})^{1/t}\;|\;\sigma(t)\in\{1,\dots,m\},t\in\{1,\dots,T\}\}$. Then the LHS of (2.11) is greater or equals to the RHS of (2.11). Now Suppose the LHS of (2.11) is larger than the RHS of (2.11), then there exist a $T^{\prime}>0$ and $A_{\sigma,T^{\prime}}$ such that
$s_{n}(A_{\sigma,T^{\prime}})<\liminf_{T\rightarrow\infty}S(T)^{1/T}.$ (2.13)
This is impossible as $S(T^{\prime})\leq s_{n}(A_{\sigma,T^{\prime}})$. Then the LHS of (2.11) must equal to the RHS of (2.11) and The RHS of (2.11) is the $\check{\rho}$ of the matrix set $\mathcal{M}$ (see, for example, Theorem 1.1 in [12]) ∎
Theorem 2.6.
The stabilizability radius satisfies the following inequality.
@@
The inequality $\tilde{\rho}\leq\check{\rho}$ is from the definition of $\tilde{\rho}$. For a given $T$, the area of the union of $S_{r,A_{\sigma,T}}$ under all $\sigma(\cdot)$ has an upper bound given by
$\left|\bigcup_{\sigma}S_{r,A_{\sigma,T}}\right|\leq\sum_{\sigma}|S_{r,A_{\sigma,T}}|\leq\sum_{\sigma}|S^{n-1}|I\left(\frac{r^{2}}{s^{2}_{n}(A_{\sigma,T})};\frac{1}{2},\frac{n-1}{2}\right)$ (2.15)
The sum above is over all possible choices of $\sigma(\cdot)$. Since I(h,1/2,(n1)/2)I(h,1/2,(n-1)/2) is the cumulative distribution function of the beta distribution, it monotonically increases for $0\leq h\leq 1$. Then
$I\left(\frac{r^{2}}{s^{2}_{n}(A_{\sigma,T})};\frac{1}{2},\frac{n-1}{2}\right)\leq I\left(\frac{r^{2}}{S(T)^{2}};\frac{1}{2},\frac{n-1}{2}\right)$ (2.16)
holds for any $\sigma(\cdot)$, which means that
$\left|\bigcup_{\sigma}S_{r,A_{\sigma,T}}\right|\leq m^{T}|S^{n-1}|I\left(\frac{r^{2}}{S(T)^{2}};\frac{1}{2},\frac{n-1}{2}\right).$ (2.17)
From the definition of $\tilde{\rho}$, for any $\lambda>\tilde{\rho}$, there exists a positive constant $c$ such that the union of pre-images, $S_{c\lambda^{T},A_{\sigma,T}}$ for all $\sigma(\cdot)$, covers the unit sphere. Hence that
$|S^{n-1}|=\left|\bigcup_{\sigma}S_{c\lambda^{T},A_{\sigma,T}}\right|\leq m^{T}|S^{n-1}|I\left(\frac{c^{2}\lambda^{2T}}{S(T)^{2}};\frac{1}{2},\frac{n-1}{2}\right).$ (2.18)
@@
Now we need the asymptotic behavior of I(h;1/2,(n1)/2)I(h;1/2,(n-1)/2) when $h$ tends to 0 to get a lower bound of $\tilde{\rho}$. Notice that
@@
for a constant $C$. Let $g(x)$ denote the integrand, then for any $n\geq 2$, as $x\rightarrow 0$, we have $g(x)\sim x^{-\frac{1}{2}}$, i.e.
@@
$(2+\epsilon)C\frac{c\lambda^{T}}{S(T)}\geq\frac{1}{m^{T}},$ (2.22)
for any $\epsilon>0$, which is
$((2+\epsilon)Cc)^{1/T}\lambda\geq\frac{S(T)^{1/T}}{m}.$ (2.23)
Let $T$ tend to infinity. By lemma 2.5, we have
$\lambda\geq\frac{\check{\rho}}{m}$ (2.24)
for any $\lambda>\tilde{\rho}$. Our lower bound is obtained by letting $\lambda$ tend to $\tilde{\rho}$. ∎
In general this is not a sharp bound, especially for large $m$. For Example 1.1, we have $\check{\rho}=1$. From Theorem 2.6, we know that $\tilde{\rho}\geq 1/2$, which is not as sharp as lower bound given by Theorem $2$ in [6] mentioned above. However Theorem 2.6 works for Example 1.2. As mentioned above, Example 1.2 has the same dynamics as Example 1.1. Then they have the same joint spectral subradius $\check{\rho}=1$, which gives the same lower bound of $\tilde{\rho}$ as expected.
From Theorem 2.6, we know that the stabilizability radius satisfies $\tilde{\rho}\geq\check{\rho}/m$. An immediate corollary of Theorem 2.6 is that when $\tilde{\rho}=0$, we have $\check{\rho}=0$. In other word, when a matrix set pointwise converges to zero quickly enough, it converges to zero uniformly.
3 Rank one singular matrices
The rank one singular matrix (in $\mathbb{R}^{n\times n}$), i.e. the singular matrix with one dimensional image, is important in the study of the stabilizability radius. Due to the one dimensional image of such a matrix, the difference between $\tilde{\rho}$ and $\check{\rho}$ is eliminated.
Proposition 3.1.
If $\mathcal{M}$ contains at least one rank one singular matrix, then $\tilde{\rho}(\mathcal{M})=\check{\rho}(\mathcal{M})$
Proof.
Let us recall that the definition of $\tilde{\rho}$ of a matrix set is given by
$\tilde{\rho}(\mathcal{M})=\sup_{x_{0}\in S^{n-1}}\inf\{\lambda\geq 0\,|\,\exists\sigma(\cdot),c>0\;s.t.\;\|x(t)\|\leq c\lambda^{t}\,\forall t\geq 0\},$ (3.1)
The definition (3.1) is slightly different but equivalent to the definition (1.4). Let $\tilde{\rho}_{x_{0}}(\mathcal{M})$ denote the infimum in (3.1), i.e.
$\tilde{\rho}(\mathcal{M})=\sup_{x_{0}\in S^{n-1}}\tilde{\rho}_{x_{0}}(\mathcal{M})$ (3.2)
Let $\mathcal{M}_{0}$ denote a $\tilde{\rho}-$minimal matrix set containing at least one singular matrix with a one-dimensional image. Let $m_{0}$ denote the cardinality of $\mathcal{M}_{0}$, where each matrix is indexed from $\{1,\dots,m_{0}\}$. Let $M_{1}$ be a singular matrix in $\mathcal{M}_{0}$ with one dimensional image, and $v_{1}$ be a normalized vector in its image, For any $x_{0}$ we have
$\|M_{1}x_{0}\|=C_{x_{0}}\|v_{1}\|,$ (3.3)
for some constant depending on $x_{0}$. Then from the definition of $\tilde{\rho}_{x_{0}}(\mathcal{M})$, we have
$\tilde{\rho}_{x_{0}}(\mathcal{M})\leq\tilde{\rho}_{M_{1}x_{0}}(\mathcal{M})=\tilde{\rho}_{v_{1}}(\mathcal{M}),$ (3.4)
for any $x_{0}\in S^{n-1}\setminus\text{Ker}(M_{1})$. Then $\tilde{\rho}(\mathcal{M})=\tilde{\rho}_{v_{1}}(\mathcal{M})$, which is independent from initial vectors. ∎
Corollary 3.2.
We have $\tilde{\rho}=\check{\rho}$ for any two dimensional matrix set with at least one singular matrix.
However, for arbitrary matrix set with singular matrices, we do not have such properties. In Example 1.2, we have $\check{\rho}=1$ but this system is pointwise stabilizable as it behaves the same as Example 1.1 (proved in [21]).
Notice that for a given $\mathcal{M}$, a subset of $\mathcal{M}$ may be sufficient to stabilize the system, leaving some matrices redundant for the purposes of stabilization. To avoid such phenomena, we have following definition.